OpenStudy (anonymous):

Solve x + y = 6 y + z = 0 x + y – z = 4

OpenStudy (mrnood):

the first equation gives you (x+y) put that into the 3rd equation to solve for z put z into the the second to solve for y then x is evident

OpenStudy (taramgrant0543664):

y=6-x Putting it into the third equation you get x+6-x-z=4 So z=-2 Putting z into the second equatio we get y+(-2)=O Therefore y=2 So now you can put it into the first equation x+2=6 X=?

OpenStudy (mrnood):

@taramgrant0543664 I think you have made a mistake in the first derivation of z (check the sign)

OpenStudy (taramgrant0543664):

Ya z should be positive two I forgot the (-) on the z itself lol